# Viscoplasticity can stabilise liquid collar motion on vertical cylinders

### v2

@@ Abstract
Liquid films coating vertical cylinders can form annular liquid collars which translate downwards under gravity. We investigate the dynamics of a thin viscoplastic liquid film coating the interior or exterior of a vertical cylindrical tube, quantifying how the yield stress modifies both the Rayleigh-Plateau instability leading to collar formation and the translation of collars down the tube. We use thin-film theory to derive an evolution equation for the layer thickness, which we solve numerically to determine the nonlinear dynamics of an initially flat layer. A flat layer is unstable to small disturbances in the free-surface height when gravity is sufficiently strong to make the fluid yield. We use matched asymptotics to derive a model describing the quasi-steady translation of a slender liquid collar when the Bond number is small. The structure of the asymptotic solution for a viscoplastic collar shares some features with the Newtonian version, but there are several novel asymptotic regions that emerge at the two ends of the collar. The global force balance, which determines the collar’s speed, is modified by a leading-order contribution from viscous drag in the collar when the liquid is viscoplastic. We use the asymptotic model to describe slow changes in collar volume when the film thicknesses ahead of, and behind, the collar are unequal. When the film thickness ahead of the collar is less than a critical value that we determine, viscoplastic collars adjust their volume and reach a steadily-translating state. This contrasts with the Newtonian problem, where the only state in which steady translation occurs is unstable to small changes in the film thickness.
@@ I Introduction
A liquid film coating the interior of a cylindrical tube, or the exterior of a cylindrical tube or fibre, can be unstable to a surface-tension-driven Rayleigh-Plateau instability, which causes an initially thin uniform coating to form annular collars [1]. Droplets may form on the exterior of cylinders or liquid plugs within the interior of tubes, if the coating is thicker [2, 3]. Jensen [4] developed a matched-asymptotics model for the gravity-driven translation of annular collars and plugs of Newtonian liquid on the interior of cylindrical tubes, when gravity is relatively weak compared to capillary effects. It was shown that the asymptotic solution has a three-region structure, with short regions matching the main body of the collar (or plug) onto the uniform film ahead of, and behind, the collar (or plug). The speed of translation was determined via a force balance between the weight of the collar or plug and the viscous drag from the two short regions at the ends. A similar asymptotic structure has been identified in related problems where slender collars translate quickly under strong gravitational forces [5, 6, 7]. Whilst the majority of studies have focused on the motion of Newtonian collars, droplets or plugs, Yu & Hinch [8] also used matched asymptotics to study the fast translation of slender collars of power-law fluid, quantifying how the translation speed can depend on the power-law index.
Motivation for studying the formation and motion of collars and plugs inside cylindrical tubes often arises from the application to modelling mucus mechanics in lung airways [9, 4, 3]. Airway mucus has a yield stress [10], which is raised in diseases such as cystic fibrosis or chronic obstructive pulmonary disease (COPD) [11]. The yield stress can alter the dynamics of mucus transport in airways [12, 13, 14, 15, 16, 17, 18], including potentially modifying the stresses exerted on the airway wall during flows of mucus; these stresses on the airway wall can lead to epithelial cell damage [19]. Understanding the impacts of yield stress on the gravity-driven flow of airway mucus provides motivation for this investigation into the effects of viscoplasticity on the formation and motion of liquid collars inside cylindrical tubes. Additional applications may be found in industrial and manufacturing processes in which complex fluids are coated on fibres or tubes, where formation of liquid collars or droplets may be disadvantageous [2, 1].
@@
The first investigation of the nonlinear evolution of a Newtonian liquid film coating the interior of a cylindrical tube was by Hammond [22], whose model was later extended by Gauglitz & Radke [23] to study thicker films which may form plugs in the tube. Models similar to that of Gauglitz & Radke [23] have since been used to study the surface-tension-driven instability of Newtonian films in various scenarios, including where air flow also drives motion in the layer [24, 25], in which surfactant is present at the interface [26, 27], or where the liquid is viscoelastic [28]. Numerical simulations using computational fluid dynamics (CFD) have also been used to study plug formation in axisymmetric Newtonian [29, 30] and non-Newtonian layers [31, 17, 18]; these simulations are able to capture both the pre-coalescence and post-coalescence phases of plug formation. Erken et al. [17] simulated plug formation using an elastoviscoplastic model for the liquid, with parameter values fitted to those from rheometry of airway mucus, and Fazla et al. [18] later extended the model to incorporate kinematic hardening, showing that interactions between the effects of the different rheological properties in the model can modify the dynamics around plug formation. Camassa, Ogrosky & Olander [32] conducted experiments in which a Newtonian liquid layer drained down the interior of a vertical tube, and compared results to a reduced-order long-wave model. Their model could capture formation of liquid collars or plugs, and it was found to predict well the critical film thicknesses for plug formation to occur. Dietze & Ruyer-Quil [33] also studied Newtonian films coating the interior of vertical cylindrical tubes in the case that the Bond number (a measure of the strength of gravitational forces compared to surface tension) is relatively large, and determined criteria for occlusion to occur via numerical simulations. Propagation and rupture of viscoplastic liquid plugs has also been studied both numerically and experimentally in the case that motion is driven by an imposed pressure gradient rather than by gravity [12, 13, 14]. In the present study, we focus on the formation and propagation of slender liquid collars at relatively low Bond numbers.
We have previously conducted theoretical studies of the Rayleigh-Plateau instability in viscoplastic liquid layers coating cylindrical tubes, using thin-film and long-wave theories, in the case that there is no gravity [15, 16]. It was found that increasing the yield stress can have a significant stabilising effect, slowing growth of the instability and reducing the size of collars that form when the layer is thin. In that problem, it was necessary to apply a finite-amplitude perturbation to the free-surface height to make the fluid initially yield and to trigger instability. We will show that, when gravity is introduced, thin-film theory predicts that a linear instability can be triggered by arbitrarily small perturbations to the free-surface height, removing the requirement for a finite-amplitude initial perturbation. Similarly to the case of roll wave formation in viscoplastic inclined plane flow [34], linear instability is possible when gravitational forces are strong enough to make the fluid yield and flow in the base state when the layer has uniform thickness.
@@
Viscoplastic thin-film theory, which will be used in this study, predicts that regions of plug-like flow develop in the liquid layer wherever the shear stress drops below the yield stress. These regions of plug-like flow typically develop adjacent to the free surface, and are termed ‘pseudo-plugs’ [35] when the velocity within them varies along the direction of the flow, reflecting the fact that the fluid is marginally yielded within pseudo-plugs, with normal stresses and shear stresses combining to exceed the yield stress [36]. Various free-surface flows of viscoplastic fluids, including flows driven by gravity, surface tension, or both, have been investigated previously using thin-film theory [36, 37, 38, 39, 40, 41, 15, 16]. Predictions from thin-film theory have been directly compared to computational fluid dynamics (CFD) simulations or experiments in a number of these papers [40, 39, 41]. In a study of steady bubble propagation through a tube filled with viscoplastic fluid, Jalaal & Balmforth [40] utilised both a thin-film model and CFD simulations, finding generally good agreement when the liquid film between the bubble and the tube wall was thin. Jalaal et al. [39] also found good agreement between thin-film predictions and CFD results for the spreading radius of viscoplastic droplets, except when the droplet shape deviated significantly from the shallow profile assumed in their thin-film theory. In their numerical simulations, they also identified regions adjacent to the free surface of the droplets resembling weakly-yielded pseudo-plugs when the droplets were sufficiently shallow, validating the flow structure predicted by the thin-film theory.
@@
In this study, we employ thin-film theory to investigate the Rayleigh-Plateau instability of a viscoplastic liquid film coating a vertical cylinder and the motion of slender liquid collars on the cylinder. We aim to quantify the effects of viscoplasticity on the dynamics of liquid collar formation and propagation. By developing a small-Bond-number model for quasi-steady collar motion, we make theoretical predictions for how the liquid’s yield stress modifies the speed, size and stability of a translating collar. The paper is organised as follows. In §II, we derive a thin-film evolution equation for the flow in the liquid layer, and briefly discuss methods for solving it numerically. We present results for the evolution of initially uniform films in §III, with linear stability analysis in §III.1 and an illustrative example of the nonlinear evolution of an initially uniform film discussed in §III.2. In §IV, we derive a model for quasi-steadily translating collars using matched asymptotics, detailing the solution in each of the asymptotic regions sequentially. In doing so, we extend and modify the approach used in previous studies of Newtonian fluid collars [4]. We present results from the asymptotic model in §V, focusing on steadily-translating collars in §V.1, where we also present comparisons between asymptotic and numerical results, and focusing on the asymptotic model’s predictions for the unsteady dynamics of translating collars presented in §V.2.
@@ II.1 Governing equations
We consider a vertical cylindrical tube of radius $a$, lined on the inside by a layer of viscoplastic liquid (figure 1). The rest of the tube is filled with an inviscid gas, in which the pressure is assumed to be uniform. We assume that flow in the liquid layer is axisymmetric, so the system can be described by coordinates $(r^{*},z^{*})$, where the tube wall lies at $r^{*}=a$, and gravity acts in the negative $z^{*}$-direction. We take the gas-liquid interface to lie at $r^{*}=a-h^{*}(z^{*},t^{*})$. We assume that there is no inertia in the liquid layer and that it is incompressible, so the flow is governed by
00 0 $=$ $\partial_{z^{*}}w^{*}+\frac{1}{r^{*}}\partial_{r^{*}}(r^{*}u^{*}),$ (1)
00 0 $=$ $-\partial_{r^{*}}p^{*}+\frac{1}{r^{*}}\partial_{r^{*}}(r^{*}\tau_{rr}^{*})+\partial_{z^{*}}\tau_{rz}^{*}-\frac{\tau^{*}_{\theta\theta}}{r^{*}},$ (2)
00 0 $=$ $-\partial_{z^{*}}p^{*}+\frac{1}{r^{*}}\partial_{r^{*}}(r^{*}\tau_{rz}^{*})+\partial_{z^{*}}\tau_{zz}^{*}-\rho g,$ (3)
@@
in a periodic domain, $0\leq z^{*}\leq\mathcal{L}^{*}$. In the second scenario, detailed in §IV, we assume that there is an isolated collar of liquid in a long tube, with the film thickness taking constant values ahead of and behind the collar, $h^{*}=h^{*-}_{\infty}$ ahead and $h^{*}=h^{*+}_{\infty}$ behind. These far-field film thicknesses, $h^{*\pm}_{\infty}$, may be much smaller than the characteristic thickness, $h^{*}_{0}$, of the collar itself.
@@ II.2 Non-dimensionalisation and thin-film scaling
00 0 $=$ $\partial_{y}u+\partial_{z}w,$ (11)
00 0 $=$ $\partial_{y}p,$ (12)
00 0 $=$ $-\partial_{z}p+\partial_{y}\tau_{yz}-B,$ (13)
@@ II.3 Numerical methods
To solve (22) numerically, we employed the regularisation introduced by Jalaal [42], which has also been used in various other thin-film viscoplastic flows [40, 39, 15, 16]. We introduce the regularisation parameter, $Y_{\mathrm{min}}$, and replace $Y$ in (22) by $\bar{Y}=\max(Y_{\mathrm{min}},Y)$. This regularisation introduces weak flow where the thin-film theory predicts no flow, but does not affect the evolution equation where $Y>Y_{\mathrm{min}}$. We use $Y_{\mathrm{min}}=10^{-6}$, after having checked that this value is small enough that results are independent of the exact value. We solve the regularised version of (22) using second-order centred finite differences to approximate the spatial derivatives, and an ODE solver in Matlab to solve the resulting set of equations through time. The number of finite-difference grid points used varied between $400$ and $1000$. Where fewer grid points were used, it was checked that results were not affected in any significant way by the exact number of grid points used. In §III.1, we consider the linear instability of a uniform layer with no regularisation of the evolution equation, and in the small-$B$ model derived in §IV, there is also no regularisation of the equations.
III Instability and time evolution of an initially uniform film
In this section, we consider solutions to (22) for which the layer initially has uniform thickness (23). First, we briefly present linear stability analysis of an initially uniform layer, before subsequently presenting an example numerical solution of (22) in a domain, $0\leq z\leq\mathcal{L}$, with periodic boundary conditions, to examine the nonlinear evolution.
III.1 Linear stability analysis
From the definition of $Y$ in (22), with the initial state given by (23), the fluid is initially flowing if and only if $B>J$, or, equivalently, $\mathcal{J}<1$, where we define

### v3

@@ Abstract
Liquid films coating vertical cylinders can form annular liquid collars which translate downwards under gravity. We investigate the dynamics of a thin viscoplastic liquid film coating the interior or exterior of a vertical cylindrical tube, quantifying how the yield stress modifies both the Rayleigh-Plateau instability leading to collar formation and the translation of collars down the tube. We use thin-film theory to derive an evolution equation for the layer thickness, which we solve numerically to examine the nonlinear dynamics. Instability and collar formation occur when gravity is sufficiently strong to make the fluid yield initially. We use matched asymptotics to derive a model describing the quasi-steady translation of a slender liquid collar when the Bond number is small. The structure of the asymptotic solution for a viscoplastic collar shares some features with the Newtonian version, but there are several novel asymptotic regions that emerge at the two ends of the collar. The global force balance, which determines the collar’s speed, is modified by a leading-order contribution from viscous drag in the collar when the liquid is viscoplastic. We use the asymptotic model to describe slow changes in collar volume when the film thicknesses ahead of, and behind, the collar are unequal. When the film thickness ahead of the collar is less than a critical value that we determine, viscoplastic collars adjust their volume and reach a steadily-translating state. This contrasts with the Newtonian problem, where the only state in which steady translation occurs is unstable to small changes in the film thickness.
@@ I Introduction
A liquid film coating the interior of a cylindrical tube, or the exterior of a cylindrical tube or fibre, can be unstable to a surface-tension-driven Rayleigh-Plateau instability, which causes an initially thin uniform coating to form annular collars [1]. Droplets may form on the exterior of cylinders or liquid plugs within the interior of tubes, if the coating is thicker [2, 3]. Motivation for studying the formation and motion of collars and plugs inside cylindrical tubes often arises from the application to modelling mucus mechanics in lung airways [4, 5, 3]. Airway mucus has a yield stress [6], which is raised in diseases such as cystic fibrosis or chronic obstructive pulmonary disease (COPD) [7]. The yield stress can alter the dynamics of mucus transport in airways [8, 9, 10, 11, 12, 13, 14], including potentially modifying the stresses exerted on the airway wall during flows of mucus; these stresses on the airway wall can lead to epithelial cell damage [15]. Understanding the impacts of yield stress on the gravity-driven flow of airway mucus provides motivation for this investigation into the effects of viscoplasticity on the formation and motion of liquid collars inside cylindrical tubes. Additional applications may be found in industrial and manufacturing processes in which complex fluids are coated on fibres or tubes, where formation of liquid collars or droplets may be disadvantageous [2, 1].
Jensen [5] developed a matched-asymptotics model for the gravity-driven translation of annular collars and plugs of Newtonian liquid on the interior of cylindrical tubes, when gravity is relatively weak compared to capillary effects. The limit of weak gravity compared to surface tension is relevant when modelling the small conducting airways in the lungs [5]. Jensen [5] showed that the asymptotic solution for a Newtonian collar (or plug) has a three-region structure, with short regions matching the main body of the collar (or plug) onto the uniform film ahead of, and behind, the collar (or plug). The speed of translation was determined via a force balance between the weight of the collar or plug and the viscous drag from the two short regions at the ends. A similar asymptotic structure has been identified in related problems where slender collars translate quickly under strong gravitational forces [16, 17, 18]. Whilst the majority of studies have focused on the motion of Newtonian collars, droplets or plugs, Yu & Hinch [19] also used matched asymptotics to study the fast translation of slender collars of power-law fluid, quantifying how the translation speed can depend on the power-law index.
@@
The first investigation of the nonlinear evolution of a Newtonian liquid film coating the interior of a cylindrical tube was by Hammond [22], whose model was later extended by Gauglitz & Radke [23] to study thicker films which may form plugs in the tube. Models similar to that of Gauglitz & Radke [23] have since been used to study the surface-tension-driven instability of Newtonian films in various scenarios, including where air flow also drives motion in the layer [24, 25], in which surfactant is present at the interface [26, 27], or where the liquid is viscoelastic [28]. Numerical simulations using computational fluid dynamics (CFD) have also been used to study plug formation in axisymmetric Newtonian [29, 30] and non-Newtonian layers [31, 13, 14]; these simulations are able to capture both the pre-coalescence and post-coalescence phases of plug formation. Erken et al. [13] simulated plug formation using an elastoviscoplastic model for the liquid, with parameter values fitted to those from rheometry of airway mucus, and Fazla et al. [14] later extended the model to incorporate kinematic hardening, showing that interactions between the effects of the different rheological properties in the model can modify the dynamics around plug formation. Camassa, Ogrosky & Olander [32] conducted experiments in which a Newtonian liquid layer drained down the interior of a vertical tube, and compared results to a reduced-order long-wave model. Their model could capture formation of liquid collars or plugs, and it was found to predict well the critical film thicknesses for plug formation to occur. Dietze & Ruyer-Quil [33] also studied Newtonian films coating the interior of vertical cylindrical tubes in the case that the Bond number (a measure of the strength of gravitational forces compared to surface tension) is relatively large, and determined criteria for occlusion to occur via numerical simulations. Propagation and rupture of viscoplastic liquid plugs has also been studied both numerically and experimentally in the case that motion is driven by an imposed pressure gradient rather than by gravity [8, 9, 10]. In the present study, we focus on the formation and propagation of slender liquid collars at relatively low Bond numbers. We have previously conducted theoretical studies of the Rayleigh-Plateau instability in viscoplastic liquid layers coating cylindrical tubes, using thin-film and long-wave theories, in the case that there is no gravity [11, 12]. It was found that increasing the yield stress can have a significant stabilising effect, slowing growth of the instability and reducing the size of collars that form when the layer is thin.
@@
Viscoplastic thin-film theory, which will be used in this study, predicts that regions of plug-like flow develop in the liquid layer wherever the shear stress drops below the yield stress. These regions of plug-like flow typically develop adjacent to the free surface, and are termed ‘pseudo-plugs’ [34] when the velocity within them varies along the direction of the flow, reflecting the fact that the fluid is marginally yielded within pseudo-plugs, with normal stresses and shear stresses combining to exceed the yield stress [35]. Various free-surface flows of viscoplastic fluids, including flows driven by gravity, surface tension, or both, have been investigated previously using thin-film theory [35, 36, 37, 38, 39, 40, 41, 11, 12, 42]. For example, Ball & Balmforth [41] used thin-film theory to study the flow of a viscoplastic film coating the interior of a rotating horizontal cylinder, predicting how liquid pools at the base of the cylinder in the case that flow is primarily driven by gravity and surface tension plays a relatively minor role. Predictions from thin-film theory have been directly compared to computational fluid dynamics (CFD) simulations or experiments in a number of studies of surface-tension-driven flows [39, 38, 40]. In a study of steady bubble propagation through a tube filled with viscoplastic fluid, Jalaal & Balmforth [39] utilised both a thin-film model and CFD simulations, finding generally good agreement when the liquid film between the bubble and the tube wall was thin. Jalaal et al. [38] also found good agreement between thin-film predictions and CFD results for the spreading radius of viscoplastic droplets, except when the droplet shape deviated significantly from the shallow profile assumed in their thin-film theory. In their numerical simulations, they also identified regions adjacent to the free surface of the droplets resembling weakly-yielded pseudo-plugs when the droplets were sufficiently shallow, validating the flow structure predicted by the thin-film theory. At the leading edge of a spreading droplet [38] or at the edge of the end-cap of a translating bubble in a tube [39], there are capillary undulations at the liquid’s free surface, which coincide with rapid changes in the sign of the shear stress and direction of flow. It is well-established that viscoplastic thin-film theory breaks down in asymptotically small regions around points where the shear stress changes sign [36], but comparisons with CFD simulations in studies such as [38] and [39] suggests the theory can still provide accurate predictions of shallow surface-tension-driven flows.
@@
In this study, we employ thin-film theory to investigate the Rayleigh-Plateau instability of a viscoplastic liquid film coating a vertical cylinder and the motion of slender liquid collars on the cylinder. We aim to quantify the effects of viscoplasticity on the dynamics of liquid collar formation and propagation. By developing a small-Bond-number model for quasi-steady collar motion, we make theoretical predictions for how the liquid’s yield stress modifies the speed, size and stability of a translating collar. The paper is organised as follows. In §II, we derive a thin-film evolution equation for the flow in the liquid layer, and briefly discuss methods for solving it numerically. We present results for the evolution of initially almost flat films in §III, with discussion of how instability can be triggered by sinusoidal free-surface perturbations in §III.1 and an illustrative example of the nonlinear evolution of a layer presented in §III.2. In §IV, we derive a model for quasi-steadily translating collars using matched asymptotics, detailing the solution in each of the asymptotic regions sequentially. In doing so, we extend and modify the approach used in previous studies of Newtonian fluid collars [5]. We present results from the asymptotic model in §V, focusing on steadily-translating collars in §V.1, where we also present comparisons between asymptotic and numerical results, and focusing on the asymptotic model’s predictions for the unsteady dynamics of translating collars presented in §V.2.
@@ II.1 Governing equations
We consider a vertical cylindrical tube of radius $a$, lined on the inside by a layer of viscoplastic liquid (figure 1a). The rest of the tube is filled with an inviscid gas, in which the pressure is assumed to be uniform. We assume that flow in the liquid layer is axisymmetric, so the system can be described by coordinates $(r^{*},z^{*})$, where the tube wall lies at $r^{*}=a$, and gravity acts in the negative $z^{*}$-direction. We take the gas-liquid interface to lie at $r^{*}=a-h^{*}(z^{*},t^{*})$. We assume that there is no inertia in the liquid layer and that it is incompressible, so the flow is governed by
$0$ $=$ $\partial_{z^{*}}w^{*}+\frac{1}{r^{*}}\partial_{r^{*}}(r^{*}u^{*}),$ (1)
$0$ $=$ $-\partial_{r^{*}}p^{*}+\frac{1}{r^{*}}\partial_{r^{*}}(r^{*}\tau_{rr}^{*})+\partial_{z^{*}}\tau_{rz}^{*}-\frac{\tau^{*}_{\theta\theta}}{r^{*}},$ (2)
$0$ $=$ $-\partial_{z^{*}}p^{*}+\frac{1}{r^{*}}\partial_{r^{*}}(r^{*}\tau_{rz}^{*})+\partial_{z^{*}}\tau_{zz}^{*}-\rho g,$ (3)
@@
in a periodic domain, $0\leq z^{*}\leq\mathcal{L}^{*}$. In the second scenario, detailed in §IV, we assume that there is an isolated collar of liquid in a long tube, with the film thickness taking a constant value, $h^{*}=h^{*-}_{\infty}$, ahead of the collar and another value, $h^{*}=h^{*+}_{\infty}$, far behind the collar. We derive a model for the slow time evolution of the collar in the limit $B\rightarrow 0$, where $B$ is a Bond number that we define below. The far-field film thicknesses, $h^{*\pm}_{\infty}$, are assumed to be much smaller than the characteristic thickness, $h^{*}_{0}$, of the collar itself. When solving for the time evolution of the isolated collar, we take $h_{0}^{*}=V^{*}_{0}/(16\pi a^{2})$, where $V_{0}^{*}$ is the initial volume of the collar, excluding the volume of the very thin films ahead of and behind the collar. The exact choice of $h_{0}^{*}$ is not of unique importance, except that $h_{0}^{*}\ll a$; rescaling $h_{0}^{*}$ by a constant $O(1)$ factor would not affect the results in §IV.
@@ II.2 Non-dimensionalisation and thin-film scaling
$0$ $=$ $\partial_{y}u+\partial_{z}w,$ (11)
$0$ $=$ $\partial_{y}p,$ (12)
$0$ $=$ $-\partial_{z}p+\partial_{y}\tau_{yz}-B,$ (13)
@@ II.3 Numerical methods
To solve (22) numerically, we employed the regularisation introduced by Jalaal [43], which has also been used in various other thin-film viscoplastic flows [39, 38, 11, 12]. We introduce the regularisation parameter, $Y_{\mathrm{min}}$, and replace $Y$ in (22) by $\bar{Y}=\max(Y_{\mathrm{min}},Y)$. This regularisation introduces weak flow where the thin-film theory predicts no flow, but does not affect the evolution equation where $Y>Y_{\mathrm{min}}$. We use $Y_{\mathrm{min}}=10^{-6}$, after having checked that this value is small enough that results are independent of the exact value. We solve the regularised version of (22) using second-order centred finite differences to approximate the spatial derivatives, and an ODE solver in Matlab to solve the resulting set of equations through time. The number of finite-difference grid points used varied between $400$ and $1000$. Where fewer grid points were used, it was checked that results were not affected in any significant way by the exact number of grid points used. In §III.1, we discuss the stability of an almost flat layer with no regularisation of the evolution equation, and in the small-$B$ model derived in §IV, there is also no regularisation of the equations.
III Instability and time evolution of a viscoplastic film
III.1 Instability of an almost flat layer
In this section, we consider the stability of an almost flat layer, a uniform film that has been subject to some perturbation in the free-surface height, $h(z,t)$. For an exactly uniform film (23), from the definition of $Y$ in (22) the fluid is flowing if and only if $B>J$, or, equivalently, $\mathcal{J}<1$, where we define